OpenStudy (anonymous):

how to factor 1 - 100x^2

OpenStudy (cwrw238):

another difference of 2 squares

OpenStudy (ajprincess):

\(1-100x^2\) \(=1^2-{10}^2x^2\) \(=1^2-(10x)^2\) Difference of squares. \(a^2-b^2=(a-b)(a+b)\) Here a=1 and b=10x

OpenStudy (anonymous):

Thanks

OpenStudy (ajprincess):

yw:)